The mean score of a group of 24 students in an educational achievement test in reading was 45 with a standard deviation of 6. At the end of the school year, the mean score on an equivalent form of the same test was 50 with a standard deviation of 5. Has the class made significant progress in reading during the year? Use alpha equal to 0.05.

1
Expert's answer
2022-01-04T10:54:22-0500

Let "X_{1}" denote the score at the beginning of the school year and "X_{2}" denote the score at the end of the school year. let us assume that "x_{1} \sim N\left(\mu_{1}, \sigma_{1}{ }^{2}\right)" and "x_{2} \sim N\left(\mu_{2}, \sigma_{2}{ }^{2}\right)" independently. Assume that "\sigma_{1}{ }^{2}=\sigma_{2}{ }^{2}=\sigma^{2}" (say)

We want to determine if the lass has made a significant progress in reading during the year. That is, the hypothesis is,

"H_{0}: \mu_{1}=\mu_{2} \quad r / s \quad H_{a}: \mu_{1}<\mu_{2}"

The appropriate teat statistic is,

"T=\frac{\bar{X}_{1}-\bar{X}_{2}}{s \sqrt{\frac{1}{n_{1}}+\frac{1}{n_{2}}}} \sim_{H_o} t_{n_{1}+n_{2}-2}"

where "\bar{X}_{1}" and "\bar{X}_{2}" are the sample mean and s denote the pooled s.d.

Given,

"\begin{aligned}

\bar{X}_{1} &=45, \bar{X}_{2}=50, n_{1}=24, n_{2}=24 \\

s^{2} &=\frac{\left(n_{1}-1\right) s_{1}^{2}+\left(n_{2}-1\right) A_{2}^{2}}{n_{1}+n_{2}-2}=\frac{23 \times 36+23 \times 25}{46} \\

&=30.5

\end{aligned}"

"\begin{aligned}

\therefore s &=5.5 \\

\therefore t &=\frac{45-50}{5.5 \sqrt{\frac{2}{24}}}=\frac{-5}{1.587}=-3.149

\end{aligned}"

where t is the observed value of "T \sim t_{46}"

We reject "H_{0}" at "\alpha" level of significance, if, "t<-t_{\alpha ; 46} \quad" ["\because" the test is left tailed ]

"\begin{aligned}

&\text { For } \alpha=0.05, \\

&-t_{\alpha ; 46}=-t_{0.05 ; 46}=-1.678 \\

&\therefore t=-3.149<-t_{0.05 ; 46}

\end{aligned}"

"\therefore" we reject "H_{0}" and conclude that there in sufficient evidence to say that the class made a significant progress in reading during the year.